# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 4.1.2 Adaptive Choice of Feedback Type
In addition to selecting queries within a given feedback type, it is also possible to actively select the feedback type itself [75]. The best choice of feedback type can depend on many factors, such as human rationality as well as task-dependent factors, some of which may change during the labeling process. [91] formalize this setting as one in which we try to select a feedback design (or feedback type) $x$ out of the space of possible designs $\mathcal{X}$ such that the expected information gain over the distribution of reward functions is maximized for the next human response. Concretely, the goal is to choose a feedback design by means of
@@
[91] find that the most informative feedback type depends on the (type-dependent) rationality of the human labeler (see Section 4.2.1). More precisely, it is shown that the most informative feedback depends on the rationality factor, e.g., while demonstrations are more informative than comparisons when the human is highly rational, comparisons should be preferred in less-rational settings. Given that this rationality might change due to factors such as fatigue or an individual labeler’s capabilities, this suggests that adaptively adjusting the feedback type during the labeling process may be worthwhile. Further study of this relationship is a promising area for future work.
@@ 4.2 Challenges of Human Labeling
This section explores the label collection process, which follows after query selection. This task intersects with several related disciplines, especially within the social sciences, as it encompasses the design of interactions to facilitate informative query responses. A prominent field in this area is psychometrics [84], which focuses on measuring psychological attributes, including preferences. Similarly, survey research [76] is dedicated to developing techniques for gathering information from individuals via surveys. Human-computer interaction plays a significant role as well, investigating the design of user interfaces tailored for preference elicitation [206]. Moreover, preference label collection is also necessary for discrete choice experiments within health economics [222], where it is used for the assessment of service values.
@@ 4.2.1 Psychology-Aware Preference Elicitation
Understanding human psychology is essential for effective preference elicitation in RLHF systems. Human decision-making is complex, often diverging from traditional rational choice models due to cognitive, social, and emotional factors. This complexity is exemplified by phenomena like fatigue, which can affect the reliability of choices based on the order of queries. This section investigates these phenomena, exploring how constructive preferences, biases, framing effects, and social interactions shape the observed choices. Recognizing and addressing these psychological underpinnings is key to developing more accurate and reliable systems. In this section, we will discuss various psychological phenomena, such as cognitive biases and response biases, and related effects (fallacies, biases, heuristics, psychological phenomena impacting decision-making processes), which may falsify labels by adding systematic bias or noise.
@@
Preference learning methods typically assume the existence of inherent, stable preferences that can be elicited through querying. Contrary to this assumption, psychological research, such as the work by [163], indicates that preferences are often constructed during the elicitation process and may vary with the method of elicitation or over time. This suggests that the feedback type not only affects elicitation’s effectiveness but also shapes preferences. Systematic biases, noise, and other psychological factors may influence observed choices, challenging the traditional models of human choice used to infer latent utilities (see Section 5.1). The elicitation method, query presentation, and context thus play a critical role in shaping measured preferences, compounded by cognitive biases and irrationalities.
@@
The influence of psychological phenomena on preference learning has been well-documented in the literature, especially within the context of explicit preference elicitation for recommender systems. For instance, [252] provide a thorough discussion of the relationship between psychology and recommender systems. Similarly, [10] review how preference construction is influenced by cognitive biases, personality traits, and emotional states in recommender systems, discussing effects like serial position, framing, anchoring, choice overload, and preference visibility. In a more specialized discussion, [176] focus on cognitive biases in the context of consumer decision-making and its interaction with recommender systems. [176] specifically address cognitive biases in consumer decision-making in interaction with recommender systems. Finally, [136] link these psychological aspects to RLHF, discussing the common practice of using synthetic instead of real human feedback for algorithm evaluation and highlighting the limitations of that approach. They further discuss challenges posed by real human feedback, many of which are related to the concepts discussed in the following paragraphs, as well as the opportunities provided by integrating psychological insights into RLHF systems.
@@
Constructive preferences are closely related to framing effects, which refer to changes in elicited preferences based on how tasks or alternatives are described, even when these descriptions are essentially equivalent. For example, presenting a choice as a loss versus a gain can lead to different decisions despite identical outcomes. Moreover, serial position effects, commonly known as primacy and recency effects, also play a significant role. These effects describe the tendency for the beginning and end of an experience to influence subjective experience disproportionately. This phenomenon becomes particularly relevant in scenarios like video choices, where the initial or concluding segments might disproportionately affect preferences. [10] discuss both of these effects in the context of recommender systems.
@@
Ordering effects pose another challenge in preference elicitation, where the sequence of queries can affect responses. [66] outline several factors contributing to these effects: institutional learning, changing preferences, and varying levels of cognitive effort. Institutional learning involves gaining familiarity with the task and feedback type, which can enhance labelers’ expertise and, consequently, the accuracy of their responses. However, due to the constructive nature of preferences, their preferences may evolve during the elicitation process, leading to changing preferences. This evolution might also be influenced by anchoring effects, where previously seen instances bias responses. Furthermore, cognitive effort levels can fluctuate due to factors like fatigue or boredom. This is closely related to choice overload, a form of fatigue from excessive choices, as discussed by [10] and bounded rationality, as explored by [51]. In such scenarios, labelers might opt out of making a choice when overwhelmed by options. Bounded rationality refers to the limitations in human decision-making capabilities, particularly when processing large amounts of information, which aligns with the concept of choice overload. To address these challenges, studies like [30] and [288] propose methods to reduce cognitive effort in responding to queries. [30] focus on posing queries that are straightforward for humans to answer, while [288] enhance the human evaluation process by presenting queries in a user-friendly format.
@@
When multiple labelers collaborate on the same task in preference elicitation, as is studied, e.g., by [16] and [65], this may lead to another set of biases if they have the opportunity to exchange information. This exchange may either be direct or indirect through observing the system’s predictions, which are based on the other labeler’s feedback. Such interactions can affect their preferences through several mechanisms, as identified by [10]: anchoring effects, transfer of emotional states, and conflict avoidance. Anchoring effects, for instance, occur when a labeler’s choices are influenced by the knowledge of others’ preferences or system predictions, a phenomenon also discussed under the term preference visibility. This bias can lead labelers to align their preferences with the anchors they are exposed to, which is a significant consideration in recommender systems. Understanding these biases is crucial for designing RLHF systems that mitigate the influence of labeler interactions on preference construction.
@@
The effects previously discussed stem from systemic biases in preference expression. In addition to these biases, choices may also be affected by noise. This is commonly discussed under the term stochastic rationality, where an agent’s behavior is rational with respect to an unobserved random state. The reward-rational implicit choice framework, as introduced by [125], addresses this by integrating a rationality factor $\beta$ into the human choice model (see Eq. 3). This factor’s impact has been further examined by [91] through synthetic experiments and user studies, demonstrating that accurately estimating this type-dependent rationality coefficient can enhance learning performance and guide feedback type selection (see Section 4.1.2). However, a practical method for estimating this factor remains a challenge. While [91] use calibration feedback with a known latent utility function for estimation, such an approach is not feasible for most tasks. In a related study, [65] investigate a scenario with multiple teachers, focusing on the agent’s ability to select the most knowledgeable or rational teacher. Therefore, developing more advanced methods to estimate this factor, along with understanding its variability due to factors like fatigue or other ordering effects, presents a vital area for future research in preference elicitation
@@
Incorporating psychological insights into the preference-learning components of RLHF systems is essential for optimizing their efficacy. A key area of focus should be research aimed at mitigating biases and harnessing cognitive aspects of preference formation. For instance, designing user interfaces that minimize framing effects and developing algorithms that account for ordering and serial positioning are crucial steps. In this realm, [181] propose a configurable user interface, called RLHF-Blender, for studying various feedback modalities and their combinations. Additionally, the study by [148] on the impact of feedback type, such as binary critiques versus action advice, on task performance and labeler satisfaction highlights the significant role of feedback type in preference elicitation. Furthermore, the work of [206] in user-interaction design underlines the importance of having an expressive feedback type to increase user engagement.
@@ 4.2.2 Importance of Researcher-Labeler Agreement
High-quality labels are important for the final policy in an RLHF process. Early work on fine-tuning language models using RLHF noticed a mismatch between the researcher’s goals and the (paid) labeler’s actual labels (researcher-labeler disagreement). [296] note that researchers agreed with each other about $60\%$ of the time (on 4-way comparisons, where random choice would result in $25\%$ agreement), while agreeing with labelers only $38\%$ or $46\%$ of the time (depending on the task). [244] attempt to reduce these disagreements by maintaining a hands-on relationship with the labelers and thereby ensuring high researcher-labeler agreement. Concretely, they provide on-boarding with detailed instructions, keep an open channel of communication between researchers and labelers, and give feedback to the labelers. They evaluate the researcher-labeler agreement and reach an agreement rate of $77\%\pm 2\%$.
@@
Due to the inherently subjective nature of the task (researchers agreed with each other in $73\%\pm 4\%$ of the cases), a much higher agreement rate cannot be expected. [200] also report the agreement rates on a different task (instruction fine-tuning instead of summarization) and find that labelers agree with each other in $72.6\pm 1.5\%$ of the time, after a screening procedure that, amongst others, selects labelers that agree with researcher labels.
@@
The importance of quality does not trump the importance of quantity, however. Indeed, [244] note that excluding low-confidence samples from the data set generally did not help with reward model training. This indicates that even though quality is important, a larger quantity is still generally better.
@@
The scale of the labeled data set required for effective training and refinement varies widely, impacting the quality of the resulting models. Studies have shown a broad range in data set sizes, from tens of labels in smaller studies [122] to hundreds in more complex scenarios [58]. Larger-scale applications may require thousands [101] or even millions of labels [2], each bringing its own challenges in ensuring label accuracy and consistency. This variability in data set size underscores the need for rigorous label quality control measures across different scales. In smaller data sets, each label carries more weight, making accuracy and precision critical. Conversely, in larger data sets, the challenge lies in maintaining consistency and mitigating systematic biases that might emerge from the sheer volume of data.
@@
Similarly, the labeling setting varies in the surveyed works, from author-provided feedback [138], over small in-person studies [135], to larger remote studies [138]. Each setting provides unique challenges to ensure high-quality labels.
@@
Various works have suggested measures to improve label quality. [111] discuss the possible failure modes of the labeling task in more detail, for example, discussing systematic biases and conflicting motivation, and propose concrete changes to the training and evaluation methodology to alleviate these. [96] suggest providing labelers with multiple natural language rules and collecting preference labels for each rule individually to improve label quality. This is related to [13], who propose to generate feedback automatically based on such a set of rules and a language model.
@@ 5 Reward Model Training
In this section, we delve deeper into the aspect of reward model learning, which we briefly touched on in Section 2.3. We will discuss various aspects associated with the topic, such as the different reward model inputs or model architectures considered hitherto or the various possibilities regarding the training process.
@@ 5.1 Human Feedback Model
The basic premise underlying the majority of approaches in RLHF is that human feedback is directly related to the reward function to be learned. To this end, the human feedback must first be captured in a sound mathematical framework that establishes the connection to the reward function. On a high level, one can break down (almost) all feedback types in Section 3.2 to a choice scenario: The human chooses one specific feedback option (label) from an available (possibly infinite) pool of possible feedback options (choice sets) [^1] This point of view goes back to the work of [125].. Here, the query that is made specifies the explicit contents of the choice set, e.g., if the query is to compare two trajectories, then the choice set consists of all possible outcomes for these two trajectories.
@@
for the probability that a human labeler returns a label $l$ from all possible candidate labels that can be extracted from a given query $q$ One could also recover the noiseless scenario if the latter probability distribution is degenerated for all possible candidate label sets.

### v2

@@ 4.1.2 Adaptive Choice of Feedback Type
In addition to selecting queries within a given feedback type, it is also possible to actively select the feedback type itself (Fitzgerald et al., 2022). The best choice of feedback type can depend on many factors, such as human rationality as well as task-dependent factors, some of which may change during the labeling process. Ghosal et al. (2023) formalize this setting as one in which we try to select a feedback design (or feedback type) $x$ out of the space of possible designs $\mathcal{X}$ such that the expected information gain over the distribution of reward functions is maximized for the next human response. Concretely, the goal is to choose a feedback design by means of
@@
Ghosal et al. (2023) find that the most informative feedback type depends on the (type-dependent) rationality of the human labeler (see Section 4.2.1). More precisely, it is shown that the most informative feedback depends on the rationality factor, e.g., while demonstrations are more informative than comparisons when the human is highly rational, comparisons should be preferred in less-rational settings. Given that this rationality might change due to factors such as fatigue or an individual labeler’s capabilities, this suggests that adaptively adjusting the feedback type during the labeling process may be worthwhile. Further study of this relationship is a promising area for future work.
@@ 4.2 Challenges of Human Labeling
This section explores the label collection process, which follows after query selection. This task intersects with several related disciplines, especially within the social sciences, as it encompasses the design of interactions to facilitate informative query responses. A prominent field in this area is psychometrics (Furr, 2021), which focuses on measuring psychological attributes, including preferences. Similarly, survey research (Fowler, 2013) is dedicated to developing techniques for gathering information from individuals via surveys. Human-computer interaction plays a significant role as well, investigating the design of user interfaces tailored for preference elicitation (Pommeranz et al., 2012). Moreover, preference label collection is also necessary for discrete choice experiments within health economics (Ryan et al., 2008), where it is used for the assessment of service values.
@@ 4.2.1 Psychology-Aware Preference Elicitation
Understanding human psychology is essential for effective preference elicitation in RLHF systems. Human decision-making is complex, often diverging from traditional rational choice models due to cognitive, social, and emotional factors. This complexity is exemplified by phenomena like fatigue, which can affect the reliability of choices based on the order of queries. This section overviews these phenomena, exploring how constructive preferences, biases, framing effects, and social interactions shape the observed choices. Recognizing and addressing these psychological underpinnings is key to developing more accurate and reliable systems. In this section, we will discuss various psychological phenomena, such as cognitive biases and response biases, and related effects (fallacies, biases, heuristics, psychological phenomena impacting decision-making processes), which may falsify labels by adding systematic bias or noise.
@@
Preference learning methods typically assume the existence of inherent, stable preferences that can be elicited through querying. Contrary to this assumption, psychological research, such as the work by Lichtenstein & Slovic (2006), indicates that preferences are often constructed during the elicitation process and may vary with the method of elicitation or over time. This suggests that the feedback type not only affects elicitation’s effectiveness but also shapes preferences. Systematic biases, noise, and other psychological factors may influence observed choices, challenging the traditional models of human choice used to infer latent utilities (see Section 5.1). The elicitation method, query presentation, and context thus play a critical role in shaping measured preferences, compounded by cognitive biases and irrationalities.
@@
The influence of psychological phenomena on preference learning has been well-documented in the literature, especially within the context of explicit preference elicitation for recommender systems. For instance, Tran et al. (2021) provide a thorough discussion of the relationship between psychology and recommender systems. Similarly, Atas et al. (2021) review how preference construction is influenced by cognitive biases, personality traits, and emotional states in recommender systems, discussing effects like serial position, framing, anchoring, choice overload, and preference visibility. In a more specialized discussion, Mandl et al. (2011) focus on cognitive biases in the context of consumer decision-making and its interaction with recommender systems. Mandl et al. (2011) specifically address cognitive biases in consumer decision-making in interaction with recommender systems. Finally, Kaufmann et al. (2023) link these psychological aspects to RLHF, discussing the common practice of using synthetic instead of real human feedback for algorithm evaluation and highlighting the limitations of that approach. They further discuss challenges posed by real human feedback, many of which are related to the concepts discussed in the following paragraphs, as well as the opportunities provided by integrating psychological insights into RLHF systems.
@@
Constructive preferences are closely related to framing effects, which refer to changes in elicited preferences based on how tasks or alternatives are described, even when these descriptions are essentially equivalent. For example, presenting a choice as a loss versus a gain can lead to different decisions despite identical outcomes. Moreover, serial position effects, commonly known as primacy and recency effects, also play a significant role. These effects describe the tendency for the beginning and end of an experience to influence subjective experience disproportionately. This phenomenon becomes particularly relevant in scenarios like video choices, where the initial or concluding segments might disproportionately affect preferences. Atas et al. (2021) discuss both of these effects in the context of recommender systems.
@@
Ordering effects pose another challenge in preference elicitation, where the sequence of queries can affect responses. Day et al. (2012) outline several factors contributing to these effects: institutional learning, changing preferences, and varying levels of cognitive effort. Institutional learning involves gaining familiarity with the task and feedback type, which can enhance labeler’s expertise and, consequently, the accuracy of their responses. However, due to the constructive nature of preferences, their preferences may evolve during the elicitation process, leading to changing preferences. This evolution might also be influenced by anchoring effects, where previously seen instances bias responses. Furthermore, cognitive effort levels can fluctuate due to factors like fatigue or boredom. This is closely related to choice overload, a form of fatigue from excessive choices, as discussed by Atas et al. (2021) and bounded rationality, as explored by Chen et al. (2013). In such scenarios, labelers might opt out of making a choice when overwhelmed by options. Bounded rationality refers to the limitations in human decision-making capabilities, particularly when processing large amounts of information, which aligns with the concept of choice overload. To address these challenges, studies like Bıyık et al. (2019) and Zhang et al. (2022) propose methods to reduce cognitive effort in responding to queries. Bıyık et al. (2019) focus on posing queries that are straightforward for humans to answer, while Zhang et al. (2022) enhance the human evaluation process by presenting queries in a user-friendly format.
@@
When multiple labelers collaborate on the same task in preference elicitation, as is studied, e.g., by Barnett et al. (2023) and Daniels-Koch & Freedman (2022), this may lead to another set of biases if they have the opportunity to exchange information. This exchange may either be direct or indirect through observing the system’s predictions, which are based on the other labeler’s feedback. Such interactions can affect their preferences through several mechanisms, as identified by Atas et al. (2021): anchoring effects, transfer of emotional states, and conflict avoidance. Anchoring effects, for instance, occur when a labeler’s choices are influenced by the knowledge of others’ preferences or system predictions, a phenomenon also discussed under the term preference visibility. This bias can lead labelers to align their preferences with the anchors they are exposed to, which is a significant consideration in recommender systems. Understanding these biases is crucial for designing RLHF systems that mitigate the influence of labeler interactions on preference construction.
@@
The effects previously discussed stem from systemic biases in preference expression. In addition to these biases, choices may also be affected by noise. This is commonly discussed under the term stochastic rationality, where an agent’s behavior is rational with respect to an unobserved random state. The reward-rational implicit choice framework, as introduced by Jeon et al. (2020), addresses this by integrating a rationality factor $\beta$ into the human choice model (see Eq. 3). This factor’s impact has been further examined by Ghosal et al. (2023) through synthetic experiments and user studies, demonstrating that accurately estimating this type-dependent rationality coefficient can enhance learning performance and guide feedback type selection (see Section 4.1.2). However, a practical method for estimating this factor remains a challenge. While Ghosal et al. (2023) use calibration feedback with a known latent utility function for estimation, such an approach is not feasible for most tasks. In a related study, Daniels-Koch & Freedman (2022) investigate a scenario with multiple teachers, focusing on the agent’s ability to select the most knowledgeable or rational teacher. Therefore, developing more advanced methods to estimate this factor, along with understanding its variability due to factors like fatigue or other ordering effects, presents a vital area for future research in preference elicitation
Finally, the quality of human feedback is biased towards factors that are easy to judge. Hosking et al. (2024) demonstrates that in the case of LLM fine-tuning, humans tend to favor assertiveness over factuality, since the latter is hard to judge without external assistance or resources. A similar phenomenon was observed in the control setting by Amodei et al. (2017), where the agent learned a behavior that looked good only from the camera angle that the human labelers had access to.
@@
Incorporating psychological insights into the preference-learning components of RLHF systems is essential for optimizing their efficacy. A key area of focus should be research aimed at mitigating biases and harnessing cognitive aspects of preference formation. For instance, designing user interfaces that minimize framing effects and developing algorithms that account for ordering and serial positioning are crucial steps. In this realm, Metz et al. (2023) and Yuan et al. (2024) each propose a configurable user interface for studying various feedback types and their combinations. Additionally, the study by Krening & Feigh (2018) on the impact of feedback type, such as binary critiques versus action advice, on task performance and labeler satisfaction highlights the significant role of feedback type in preference elicitation. Furthermore, the work of Pommeranz et al. (2012) in user-interaction design underlines the importance of having an expressive feedback type to increase user engagement.
@@ 4.2.2 Importance of Researcher-Labeler Agreement
High-quality labels are important for the final policy in an RLHF process. Early work on fine-tuning language models using RLHF noticed a mismatch between the researcher’s goals and the (paid) labeler’s actual labels (researcher-labeler disagreement). Ziegler et al. (2020) note that researchers agreed with each other about $60\%$ of the time (on 4-way comparisons, where random choice would result in $25\%$ agreement), while agreeing with labelers only $38\%$ or $46\%$ of the time (depending on the task). Stiennon et al. (2020) attempt to reduce these disagreements by maintaining a hands-on relationship with the labelers and thereby ensuring high researcher-labeler agreement. Concretely, they provide on-boarding with detailed instructions, keep an open channel of communication between researchers and labelers, and give feedback to the labelers. They evaluate the researcher-labeler agreement and reach an agreement rate of $77\%\pm 2\%$.
@@
Perfect labels are often impossible due to the inherently subjective nature of the task. Returning to the example given by Stiennon et al. (2020), different researchers agreed with each other in only $73\%\pm 4\%$ of the cases. Ouyang et al. (2022) also report the agreement rates on a different task (instruction fine-tuning instead of summarization) and find that labelers agree with each other in $72.6\pm 1.5\%$ of the time, after a screening procedure that, amongst others, selects labelers that agree with researcher labels. Preferences can additionally be inconsistent between feedback types, as demonstrated by the finding of Bansal et al. (2024), which shows that preferences inferred from ratings and rankings significantly disagree for both human and AI annotators.
@@
The importance of quality does not trump the importance of quantity, however. Indeed, Stiennon et al. (2020) note that excluding low-confidence samples from the data set generally did not help with reward model training. This indicates that even though quality is important, a larger quantity is still generally better.
@@
The scale of the labeled data set required for effective training and refinement varies widely, impacting the quality of the resulting models. Studies have shown a broad range in data set sizes, from tens of labels in smaller studies (Jain et al., 2015) to hundreds in more complex scenarios (Christiano et al., 2017). Larger-scale applications may require thousands (Guan et al., 2021) or even millions of labels (Abramson et al., 2022), each bringing its own challenges in ensuring label accuracy and consistency. This variability in data set size underscores the need for rigorous label quality control measures across different scales. In smaller data sets, each label carries more weight, making accuracy and precision critical. Conversely, in larger data sets, the challenge lies in maintaining consistency and mitigating systematic biases that might emerge from the sheer volume of data.
@@
Similarly, the labeling setting varies in the surveyed works, from author-provided feedback (Kim et al., 2023), over small in-person studies (Katz et al., 2021), to larger remote studies (Kim et al., 2023). Each setting provides unique challenges to ensure high-quality labels.
@@
Various works have suggested measures to improve label quality. Hagendorff & Fabi (2022) discuss the possible failure modes of the labeling task in more detail, for example, discussing systematic biases and conflicting motivation, and propose concrete changes to the training and evaluation methodology to alleviate these. Glaese et al. (2022) suggest providing labelers with multiple natural language rules and collecting preference labels for each rule individually to improve label quality. This is related to Bai et al. (2022b), who propose to generate feedback automatically based on such a set of rules and a language model.
@@ 5 Reward Model Training
[…]
Figure 4: Highlighting the components of the RLHF loop discussed in this section.
In this section, we delve deeper into the process of reward model learning, which we briefly touched on in Section 2.3. We will discuss various aspects associated with this topic, namely the various human feedback models, utility (i.e., reward) learning per se, different reward model inputs, and how to increase feedback efficiency.
@@ 5.1 Human Feedback Model
The basic premise underlying the majority of approaches in RLHF is that human feedback is directly related to the reward function to be learned. To this end, the human feedback must first be captured in a sound mathematical framework that establishes the connection to the reward function. On a high level, one can break down (almost) all feedback types in Section 3.2 to a choice scenario: The human chooses one specific feedback option (label) from an available (possibly infinite) pool of possible feedback options (choice sets) [^2] This point of view goes back to the work of Jeon et al. (2020).. Here, the query that is made specifies the explicit contents of the choice set, e.g., if the query is to compare two trajectories, then the choice set consists of all possible outcomes for these two trajectories.
@@
for the probability that a human labeler returns a label $l$ from all possible candidate labels that can be extracted from a given query $q.$ We explain next how this probability can be modeled and discuss various related modeling questions (e.g., human rationality, multiple humans, or Markov assumption). One could also recover the noiseless scenario if the latter probability distribution is degenerated for all possible candidate label sets.